Two high school seniors find themselves with much time and nothing to do. They decide to devise a computer game in which a cursor moves randomly along the x axis between –1 and 1. The game they play is to stop the cursor at random points between –1 and 1 without looking at the screen. Such a game results in a uniform distribution over the interval between –1 and 1. Since the area under the density curve (here a flat line) equals 1, the height of the density curve is: 1. 0.5. 2. 0.25.
Answer #1

here as we know area under the density curve

therefore height *width =1

height*(1-(-1))=1

height*2 =1

height =1/2 =0.5